40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 169     Entry time: Mon Oct 14 13:40:16 2013
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: H1 OMC Optical testing 

Since the middle of September, the optical tests of H1 OMC were took place.
Here is summary of the progress.

TEST1: FSR/FINESSE measurement before applying First Contact
TEST2: Power budget

MIrror cleaning with First Contact

TEST3: FSR/FINESSE measurement after First Contact application
TEST4: Power budget

TEST5: N/A

TEST6: HOM measurement @PZT V=0
TEST7: HOM measurement @PZT V=0-200

TEST8: DC response of the PZT
TEST9: AC response of the PZT

TEST10: PD/QPD alignment / output check

 

 

ELOG V3.1.3-